OpenStudy (anonymous):

1. Give an example of the commutative property of multiplication. 2.Give an example of why division is not commutative.?? I just didn't see anything really about this in the book so... I just don't know...

OpenStudy (turingtest):

similar to what you just did you are to decide if\[\frac ab=\frac ba\]

Parth (parthkohli):

\[3 \times 4 = 4 \times 3 \]The above is commutative property of multiplication.

Parth (parthkohli):

The best example I give to my students is this one:\[ {0 \over n } \ne {n \over 0}\]

Parth (parthkohli):

Where \(0 \over n\) is 0(n \(\ne\) 0), \(n \over 0\) is undefined.

mathslover (mathslover):

\[\large{\frac{a}{b}=\frac{b}{a}}\] \[\large{ab=ba}\] these are our points... now: \[\large{ab = ba}\] \[\large{a\times b = b\times a}\] \[\large{\frac{ab}{a}=\frac{ba}{a}}\] \[\large{\frac{\cancel{a}^b}{\cancel{a}^1}=\frac{b\cancel{a}^1}{\cancel{a}^1}}\] \[\large{b = b}\]

Parth (parthkohli):

We also know that 0 is NOT undefined.

OpenStudy (anonymous):

Ok I got the multiplying part but I am still confused on the divison

OpenStudy (anonymous):

would the division be 20/5 = 5/20?

OpenStudy (anonymous):

Yeah, as far as I can tell. Obviously 20/5 doesn't equal 5/20. The order of the terms matters in division, so it is not commutative. In general:\[\frac{a}{b}\neq \frac{b}{a}\]

OpenStudy (anonymous):

Ok thanks for the help.

OpenStudy (anonymous):

no prob